OpenStudy (anonymous):

which of the following is a factor or the trinomial below x^2 + x -12 A.(x-3) B. (x+3) C.(x+6) D.(x-4)

OpenStudy (xapproachesinfinity):

What are the two numbers if you multiply them you get 12 and if you add them you get 1. Think about factors of 12

OpenStudy (anonymous):

alight where do i go from there ?

OpenStudy (xapproachesinfinity):

Do you know the two numbers or not

OpenStudy (anonymous):

two numbers that you get to 12 ?

OpenStudy (xapproachesinfinity):

Yes if you multiply them you get 12

OpenStudy (anonymous):

2 and 6 ?

OpenStudy (xapproachesinfinity):

True but those two if you add them you don't get 1. What else

OpenStudy (anonymous):

okay you have to add it and get one thats what i wasnt getting

OpenStudy (xapproachesinfinity):

Give me a moment I will open my laptop ok

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

Alright! know I can Explain to you nicely

OpenStudy (xapproachesinfinity):

(x+a)(x+b)=x^2+(a+b)x+ab. we're doing this factorization

OpenStudy (xapproachesinfinity):

a.b= -12, a+b= 1 do you agree

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

Good! now what we need to do is to find two number that will give us that -12 and if we add them they give us 1. you sad 2 and 6. let's check

OpenStudy (anonymous):

thats not going to work ik already

OpenStudy (xapproachesinfinity):

2x6=12, 2+6= 8, let's check negative case -2x6=-12 , -2 +6=4

OpenStudy (xapproachesinfinity):

yeah i know I'm just going to show you how it works

OpenStudy (xapproachesinfinity):

now let's think about another two numbers that multiply to 12

OpenStudy (xapproachesinfinity):

can you tell me?

OpenStudy (anonymous):

cant think of any that would add to one

OpenStudy (xapproachesinfinity):

just for the moment think about any two that multiply to 12 okay do you know what are the factor of 12?

OpenStudy (xapproachesinfinity):

factor of 12 are the numbers that divide 12? you know this right

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

{1,2,3,4,6,12} Do you agree that this is the factors of 12?

OpenStudy (anonymous):

yeah

OpenStudy (xapproachesinfinity):

we tried 2 and 6, can we try now 3 nd 4?

OpenStudy (xapproachesinfinity):

think about negative sign too not just 3 and 4 it could be -3,4 or -4,3. well actually you need a negative there since we need -12

OpenStudy (xapproachesinfinity):

do you agree

OpenStudy (anonymous):

yeah we need to use -3 ,4

OpenStudy (xapproachesinfinity):

ok can you try that if it works?

OpenStudy (anonymous):

it does you get one i did it

OpenStudy (xapproachesinfinity):

take your time and do it neatly

OpenStudy (xapproachesinfinity):

Good job!

OpenStudy (xapproachesinfinity):

So now lets put it in this since a=-3 and b=4 we say that (x-3)(x+4)=x^2+x-12 do you agree

OpenStudy (xapproachesinfinity):

that's what we found right?

OpenStudy (anonymous):

yep

OpenStudy (xapproachesinfinity):

So x^2+x-12 is factorized to (x-3)(x+4) can we say the answer to the question is A: (x-3)

OpenStudy (xapproachesinfinity):

because if we multiply (x-3) with (x+4) we get the original trinomial

OpenStudy (anonymous):

ohh i get it more than i did thank you

OpenStudy (xapproachesinfinity):

Great job, you did here! you see you can do this. practice more of this and you will master it. And pleasure to help you!

OpenStudy (xapproachesinfinity):

Just follow the same technique, it will always work for you

OpenStudy (anonymous):

will do

OpenStudy (xapproachesinfinity):

Okay good!